OpenStudy (anonymous):

simplify: (2sqrt-50)(1/8*sqrt-2)

OpenStudy (kropot72):

Is this the problem to be solved? \[(2\times \sqrt{-50})\div(8\times \sqrt{-2})\]

OpenStudy (anonymous):

\[(2\]